OpenStudy (anonymous):

Find the shortest (straight line) distance from the positive x-axis to the positive y-axis passing through the point (8; 1).

OpenStudy (anonymous):

Which ordered pair is a solution of the equation 4x − 2y = 16?

OpenStudy (anonymous):

Looking for the formulas to start, though of \[(x + 8)^2 + (y + 1)^2 = D^2\] would be a start

OpenStudy (anonymous):

Another possible formula may be the proportion of the area of the larger triangle formed vs the area of the smaller (4 square units)???